OpenStudy (winsteria):

SOMEONE PLEASE HELP ME IF YOU DO I'LL GIVE YOU A MEDAL AND FAN What is ~ 3 7 --- ÷ --- 8 12 A) 7 ~ --- 32 B) 9 ~ --- 14 C) 14 ~ --- 9 D) 32 ~ --- 7 (IMAGINE THE ~ SIGN AS A NEGATIVE SIGN)

OpenStudy (winsteria):

Reload it if there's question marks sorry :/

OpenStudy (anonymous):

\[(-3/8) \div (7/12) = (-3/8)*(12/7) \] \[(-3*12)/(8*7) = -36/56\] now bring -36/56 to simplest form

OpenStudy (nurali):

-3/8/7/12 -3/8*12/7 -36/56

OpenStudy (winsteria):

Thank you both so much

OpenStudy (winsteria):

so is it B or...? @mony01

OpenStudy (mony01):

yes its B

OpenStudy (winsteria):

thank youuu